ABELIAN SOLITONS

RON Y. DONAGI EMMA PREVIATO

Abstract. We describe a new algebraically completely integrable system, whose integral manifolds are co-elliptic subvarieties of Jacobian varieties. This is a multiperiodic extension of the Krichever-Treibich-Verdier system, which consists of elliptic solitons.